OpenStudy (anonymous):

How to graph 2x + 4y = 4 and -2y = x - 2

OpenStudy (anonymous):

make it into y=mx+b format

OpenStudy (anonymous):

\[2x+4y=4\] solve for Y

OpenStudy (anonymous):

4y=-2x+4

OpenStudy (anonymous):

then divide by 4 so y=-1/2(x) +1

OpenStudy (anonymous):

for -2y=x-2 just divide by 2

OpenStudy (anonymous):

so

OpenStudy (ryan123345):